# Gated Integration of Low-Rank Adaptation for Continual Learning of Language Models

### v1

@@ B.5.2 Computation of Trainable Parameters in T5-XL
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 1024}$, resulting in is $4096*r*144+r*1024*144=737280r$ trainable parameters. When $r=4$, O-LoRA has $737280*4=2949120=2.95$M trainable parameters. InfLoRA only updates $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$, resulting in $4096*r*144=589824r$ trainable parameters. When $r=4$, InfLoRA has $589824*4=2359296=2.36$M trainable parameters.
@@
GainLoRA introduces the same new gating module $g_{t}(\cdot)$ as in T5-Large, with parameters $\bm{G}_{t,1}\in\mathbb{R}^{100\times 1024}$, $\bm{G}_{t,2}\in\mathbb{R}^{1024\times 100}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 1024}$. Thus, the total number of trainable parameters for GainLoRA (O-LoRA) is $2949120+1024*100+1024*100+1024=3154944=3.15$M. The total number of trainable parameters in GainLoRA (InfLoRA) is $2359296+1024*100+1024*100+1024=2565120=2.57$M.
B.5.3 Computation of Trainable Parameters in Llama-2-7B
@@
In Llama-2-7B, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{4096\times 4096}$. The model contains 32 self-attention modules, resulting in $32*2=64$ pre-trained weights that incorporate the LoRA architecture.
@@
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$, resulting in $4096*r*64+r*4096*64=524288r$ trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is $524288*4=2097152=2.10$M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$, resulting in $4096*r*64=262144r$ trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is $262144*4=1048576=1.05$M.
@@
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 4096}$, $\bm{G}_{t,2}\in\mathbb{R}^{4096\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 4096}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is $2097152+4096*50+4096*50+4096=2510848=2.51$M. The number of trainable parameters in GainLoRA (InfLoRA) is $1048576+4096*50+4096*50+4096=1462272=1.46$M.
B.5.4 Computation of Trainable Parameters in Llama-2-13B
@@
In Llama-2-13B, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{5120\times 5120}$. The model contains 40 self-attention modules, resulting in $40*2=80$ pre-trained weights that incorporate the LoRA architecture.
@@
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{5120\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 5120}$, resulting in $5120*r*80+r*5120*80=819200r$ trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is $819200*4=3276800=3.28$M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{5120\times r}$, resulting in $5120*r*80=409600r$ trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is $409600*4=1638400=1.64$M.
@@
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 5120}$, $\bm{G}_{t,2}\in\mathbb{R}^{5120\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 5120}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is $3276800+5120*50+5120*50+5120=3793920=3.79$M. The number of trainable parameters in GainLoRA (InfLoRA) is $1638400+5120*50+5120*50+5120=2155520=2.16$M.
@@ C.1 Discussing Computational Costs Introduced by GainLoRA
Existing methods, such as O-LoRA and InfLoRA, adopt the expandable LoRA architecture shown in Figure 1 (a) and fix the integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1, allowing the model to merge the expanded LoRA branches into the pre-trained matrix at inference time, thereby avoiding additional computational costs. However, when using our gating module to integrate different LoRA branches, the LoRA branches cannot be merged into the pre-trained matrix at inference time, which introduces additional computational costs. Nevertheless, we demonstrate that these computational costs are minimal compared to the computational cost of the original language models (LMs).
Table 8 compares the floating-point operations (FLOPs) and multiply-add operations (MACs) during inference for different models with and without GainLoRA. The computation of FLOPs and MACs follows the existing project calflops ye2023calflops . Here, “Original” denotes the original LMs without any LoRA adaptation. Methods such as O-LoRA and InfLoRA avoid additional computational costs by merging their LoRA branches into the original weights during inference, resulting in FLOPs and MACs identical to the original LMs. Despite introducing additional FLOPs and MACs compared to the original LMs, GainLoRA maintains minimal computational overhead relative to the original LMs.
@@ C.2 Additional Computation Introduced by Subspace Construction
The memory and computational overhead of subspace construction in GainLoRA is minimal due to the small size of the gating module (only 3 layers, see B.4). We provide detailed analyses below.
@@
GPM requires performing SVD on matrix $H_{l}H_{l}^{T}\in\mathbb{R}^{d_{l}\times d_{l}}$, where $H_{l}$ is the feature matrix in the $l$-th layer of gating module. Based on existing conclusions trefethen2022numerical , the FLOPs for SVD on $HH^{T}$ is less than $10d_{l}^{3}$. For T5-Large ($d_{1}=d_{3}=1024$ and $d_{2}=100$), this results in $10*(2*1024^{3}+100^{3})<30GFLOPs$, which is negligible compared to a single forward pass with sequence length 128 (see Table 8). Similar calculations give the same conclusion for T5-XL, Llama-2-7B, and Llama-2-13B.
C.3 Results with standard deviation
Table 9, Table 10 and Table 11 report the results with standard deviation.
Table 9: Results with standard deviation on different task sequences using T5-large model.
Method Order 1 Order 2 Order 3 Order 4
[…]
MIGU+FT du2024unlocking - - - - 71.30$\pm$1.85 11.39$\pm$1.92 69.05$\pm$0.71 14.06$\pm$0.86
SeqLoRA 7.30$\pm$1.12 47.60$\pm$0.94 7.03$\pm$0.49 47.97$\pm$0.07 49.46$\pm$2.42 27.60$\pm$4.09 33.81$\pm$0.01 45.53$\pm$1.60
IncLoRA hu2021lora 12.33$\pm$0.56 41.93$\pm$0.17 16.65$\pm$0.91 36.56$\pm$1.30 61.19$\pm$0.85 13.63$\pm$1.27 62.46$\pm$0.34 15.92$\pm$0.46
C-LoRA smithcontinual 22.69$\pm$0.01 24.25$\pm$0.90 32.81$\pm$0.64 11.60$\pm$0.23 66.83$\pm$0.56 8.64$\pm$0.32 61.86$\pm$1.77 14.18$\pm$1.50
O-LoRA wang2023orthogonal 26.37$\pm$2.27 19.15$\pm$2.15 32.83$\pm$0.25 11.99$\pm$0.38 70.98$\pm$1.74 3.69$\pm$0.53 71.21$\pm$0.33 4.03$\pm$1.00
GainLoRA (O-LoRA) 47.84$\pm$0.16 2.26$\pm$0.06 46.84$\pm$0.11 2.91$\pm$0.13 73.37$\pm$0.01 3.02$\pm$0.81 76.01$\pm$0.49 2.49$\pm$0.12
InfLoRA liang2024inflora 39.78$\pm$0.57 7.64$\pm$0.54 39.57$\pm$0.94 8.93$\pm$0.37 75.15$\pm$0.06 4.19$\pm$0.13 75.79$\pm$0.56 3.47$\pm$0.45
GainLoRA (InfLoRA) 46.21$\pm$0.05 2.40$\pm$0.24 46.44$\pm$0.41 2.61$\pm$0.25 78.01$\pm$0.26 0.77$\pm$0.01 77.54$\pm$0.23 1.25$\pm$0.10
Method Order 1 Order 2 Order 3 Order 4
[…]
O-LoRA wang2023orthogonal 36.50$\pm$4.29 11.42$\pm$5.30 40.64$\pm$1.09 6.37$\pm$0.66 73.77$\pm$1.14 2.70$\pm$0.54 76.19$\pm$0.49 3.56$\pm$0.40
GainLoRA (O-LoRA) 50.10$\pm$0.22 3.21$\pm$0.32 49.86$\pm$0.06 3.04$\pm$0.13 78.41$\pm$0.50 2.59$\pm$0.56 77.21$\pm$0.19 3.30$\pm$0.34
InfLoRA liang2024inflora 45.61$\pm$1.28 5.60$\pm$1.35 45.85$\pm$0.10 5.10$\pm$0.32 80.22$\pm$0.04 2.09$\pm$0.11 79.43$\pm$0.03 1.71$\pm$0.09
GainLoRA (InfLoRA) 50.06$\pm$0.11 1.86$\pm$0.28 50.26$\pm$0.14 2.64$\pm$0.41 81.22$\pm$0.11 0.58$\pm$0.01 80.30$\pm$0.11 0.75$\pm$0.15
Table 11: The overall results on different task sequences with Llama-2-7B and Llama-2-13B.
Llama-2-7B Llama-2-13B
Method Order 1 Order 2 Order 1 Order 2
[…]
O-LoRA wang2023orthogonal 39.37$\pm$0.24 15.84$\pm$0.58 37.55$\pm$0.70 20.23$\pm$0.20 43.92$\pm$0.42 14.15$\pm$0.35 40.05$\pm$0.46 19.53$\pm$0.50
GainLoRA (O-LoRA) 51.10$\pm$0.91 4.96$\pm$0.56 51.14$\pm$1.01 5.57$\pm$0.65 52.47$\pm$0.24 4.78$\pm$0.27 51.68$\pm$0.63 5.86$\pm$0.44
InfLoRA liang2024inflora 42.93$\pm$0.77 11.23$\pm$0.24 39.94$\pm$0.30 15.00$\pm$0.51 43.64$\pm$0.02 14.85$\pm$0.31 45.74$\pm$0.81 10.61$\pm$0.09
GainLoRA (InfLoRA) 51.27$\pm$0.01 2.84$\pm$0.11 50.17$\pm$0.32 4.71$\pm$0.22 53.64$\pm$0.81 2.87$\pm$0.05 52.46$\pm$0.50 4.90$\pm$0.30
C.4 Varying the Architecture of Gating Module
C.4.1 Varying Function $f(\cdot)$ in Gating Module
@@
GainLoRA (InfLoRA) $\left(f(b)=|2{\rm sigmoid}(b)-1|\right)$ 46.21$\pm$0.05 2.40$\pm$0.24 46.44$\pm$0.41 2.61$\pm$0.25
GainLoRA (InfLoRA) $\left(f(b)={\rm min}\{|b|,1\}\right)$ 45.05$\pm$0.32 2.07$\pm$0.24 45.00$\pm$0.20 1.74$\pm$0.44
GainLoRA (InfLoRA) $\left(f(b)=|\sin(\frac{\pi b}{2})|\right)$ 47.48$\pm$0.03 1.21$\pm$0.52 45.03$\pm$0.67 2.37$\pm$0.34
InfLoRA 39.78$\pm$0.57 7.64$\pm$0.54 39.57$\pm$0.94 8.93$\pm$0.37
GainLoRA (O-LoRA) $\left(f(b)=|2{\rm sigmoid}(b)-1|\right)$ 47.84$\pm$0.16 2.26$\pm$0.06 46.84$\pm$0.11 2.91$\pm$0.13
GainLoRA (O-LoRA) $\left(f(b)={\rm min}\{|b|,1\}\right)$ 49.62$\pm$0.57 2.83$\pm$0.73 48.62$\pm$0.47 3.74$\pm$0.02
GainLoRA (O-LoRA) $\left(f(b)=|\sin(\frac{\pi b}{2})|\right)$ 48.49$\pm$0.92 3.84$\pm$0.54 47.20$\pm$0.85 4.69$\pm$0.65
O-LoRA 26.37$\pm$2.27 19.15$\pm$2.15 32.83$\pm$0.25 11.99$\pm$0.38
C.4.2 Varying the Shapes of Weights in Gating Module
@@
C.4.3 Varying the Layers of Gating Module

### v2

@@ B.4.2 Computation of Trainable Parameters in T5-XL
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 1024}$, resulting in is 4096r144+r1024144=737280r4096*r*144+r*1024*144=737280r trainable parameters. When $r=4$, O-LoRA has 7372804=2949120=2.95737280*4=2949120=2.95M trainable parameters. InfLoRA only updates $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$, resulting in 4096r144=589824r4096*r*144=589824r trainable parameters. When $r=4$, InfLoRA has 5898244=2359296=2.36589824*4=2359296=2.36M trainable parameters.
@@
GainLoRA introduces the same new gating module $g_{t}(\cdot)$ as in T5-Large, with parameters $\bm{G}_{t,1}\in\mathbb{R}^{100\times 1024}$, $\bm{G}_{t,2}\in\mathbb{R}^{1024\times 100}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 1024}$. Thus, the total number of trainable parameters for GainLoRA (O-LoRA) is 2949120+1024100+1024100+1024=3154944=3.152949120+1024*100+1024*100+1024=3154944=3.15M. The total number of trainable parameters in GainLoRA (InfLoRA) is 2359296+1024100+1024100+1024=2565120=2.572359296+1024*100+1024*100+1024=2565120=2.57M.
B.4.3 Computation of Trainable Parameters in Llama-2-7B
@@
In Llama-2-7B, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{4096\times 4096}$. The model contains 32 self-attention modules, resulting in 322=6432*2=64 pre-trained weights that incorporate the LoRA architecture.
@@
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$, resulting in 4096r64+r409664=524288r4096*r*64+r*4096*64=524288r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 5242884=2097152=2.10524288*4=2097152=2.10M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$, resulting in 4096r64=262144r4096*r*64=262144r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 2621444=1048576=1.05262144*4=1048576=1.05M.
@@
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 4096}$, $\bm{G}_{t,2}\in\mathbb{R}^{4096\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 4096}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 2097152+409650+409650+4096=2510848=2.512097152+4096*50+4096*50+4096=2510848=2.51M. The number of trainable parameters in GainLoRA (InfLoRA) is 1048576+409650+409650+4096=1462272=1.461048576+4096*50+4096*50+4096=1462272=1.46M.
B.4.4 Computation of Trainable Parameters in Llama-2-13B
@@
In Llama-2-13B, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{5120\times 5120}$. The model contains 40 self-attention modules, resulting in 402=8040*2=80 pre-trained weights that incorporate the LoRA architecture.
@@
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{5120\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 5120}$, resulting in 5120r80+r512080=819200r5120*r*80+r*5120*80=819200r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 8192004=3276800=3.28819200*4=3276800=3.28M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{5120\times r}$, resulting in 5120r80=409600r5120*r*80=409600r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 4096004=1638400=1.64409600*4=1638400=1.64M.
@@
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 5120}$, $\bm{G}_{t,2}\in\mathbb{R}^{5120\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 5120}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 3276800+512050+512050+5120=3793920=3.793276800+5120*50+5120*50+5120=3793920=3.79M. The number of trainable parameters in GainLoRA (InfLoRA) is 1638400+512050+512050+5120=2155520=2.161638400+5120*50+5120*50+5120=2155520=2.16M.
B.4.5 Computation of Trainable Parameters in Llama-3-8B
In Llama-3-8B, the projection weights for the query and value components have shapes $\bm{W}_{q}\in\mathbb{R}^{4096\times 4096},\bm{W}_{v}\in\mathbb{R}^{4096\times 1024}$. The model contains 40 self-attention modules, resulting in 322=6432*2=64 pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$ for query and $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$ for value, resulting in 1024r32+4096r32+r409664=425984r1024*r*32+4096*r*32+r*4096*64=425984r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 4259844=1703936=1.70425984*4=1703936=1.70M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ for query and $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ for value, resulting in 4096r32+1024r32=163840r4096*r*32+1024*r*32=163840r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 1638404=655360=0.66163840*4=655360=0.66M.
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 4096}$, $\bm{G}_{t,2}\in\mathbb{R}^{4096\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 4096}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 1703936+409650+409650+4096=2117632=2.121703936+4096*50+4096*50+4096=2117632=2.12M. The number of trainable parameters in GainLoRA (InfLoRA) is 655360+409650+409650+4096=1069056=1.07655360+4096*50+4096*50+4096=1069056=1.07M.
@@ B.4.5 Computation of Trainable Parameters in Llama-3-8B
Llama-3-8B Original (1,128) 1929.86 964.89
GainLoRA (O-LoRA) (1,128) 1930.49 965.21
GainLoRA (InfLoRA) (1,128) 1930.49 965.21
@@ C.1 Discussing Computational Costs Introduced by GainLoRA
Existing methods, such as O-LoRA and InfLoRA, adopt the expandable LoRA architecture shown in Figure 1 (a) and fix the integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1, allowing the model to merge the expanded LoRA branches into the pre-trained matrix at inference time, thereby avoiding additional computational costs. However, when using our gating module to integrate different LoRA branches, the LoRA branches cannot be merged into the pre-trained matrix at inference time, which introduces additional computational costs. Nevertheless, we demonstrate that these computational costs are minimal compared to the computational cost of the original large language models (LLMs).
Table 8 compares the floating-point operations (FLOPs) and multiply-add operations (MACs) during inference for different models with and without GainLoRA. The computation of FLOPs and MACs follows the existing project calflops [69]. Here, “Original” denotes the original LLMs without any LoRA adaptation. Methods such as O-LoRA and InfLoRA avoid additional computational costs by merging their LoRA branches into the original weights during inference, resulting in FLOPs and MACs identical to the original LLMs. Despite introducing additional FLOPs and MACs compared to the original LLMs, GainLoRA maintains minimal computational overhead relative to the original LLMs.
@@ C.2 Additional Computation Introduced by Subspace Construction
The memory and computational overhead of subspace construction in GainLoRA is minimal due to the small size of the gating module (only 3 layers, see B.3). We provide detailed analyses below.
@@
GPM requires performing SVD on matrix $H_{l}H_{l}^{T}\in\mathbb{R}^{d_{l}\times d_{l}}$, where $H_{l}$ is the feature matrix in the $l$-th layer of gating module. Based on existing conclusions [55], the FLOPs for SVD on $HH^{T}$ is less than $10d_{l}^{3}$. For T5-Large ($d_{1}=d_{3}=1024$ and $d_{2}=100$), this results in $10*(2*1024^{3}+100^{3})<30GFLOPs$, which is negligible compared to a single forward pass with sequence length 128 (see Table 8). Similar calculations give the same conclusion for T5-XL, Llama-2-7B, Llama-2-13B and Llama-3-8B.
C.3 Varying the Architecture of Gating Module
C.3.1 Varying Function $f(\cdot)$ in Gating Module
@@
GainLoRA (InfLoRA) $\left(f(b)=|2{\rm sigmoid}(b)-1|\right)$ 46.21 2.40 46.44 2.61
GainLoRA (InfLoRA) $\left(f(b)={\rm min}\{|b|,1\}\right)$ 45.05 2.07 45.00 1.74
GainLoRA (InfLoRA) $\left(f(b)=|\sin(\frac{\pi b}{2})|\right)$ 47.48 1.21 45.03 2.37
InfLoRA 39.78 7.64 39.57 8.93
GainLoRA (O-LoRA) $\left(f(b)=|2{\rm sigmoid}(b)-1|\right)$ 47.84 2.26 46.84 2.91
GainLoRA (O-LoRA) $\left(f(b)={\rm min}\{|b|,1\}\right)$ 49.62 2.83 48.62 3.74
GainLoRA (O-LoRA) $\left(f(b)=|\sin(\frac{\pi b}{2})|\right)$ 48.49 3.84 47.20 4.69
O-LoRA 26.37 19.15 32.83 11.99
C.3.2 Varying the Shapes of Weights in Gating Module
@@
C.3.3 Varying the Layers of Gating Module